# Electron transport in disordered insulating lattice under nonlinear electric field

### v2

@@ Abstract
Transport in disordered systems often occurs via the variable range hopping (VRH) in the dilute carrier density limit, where electrons hop between randomly distributed localized levels. We study the nonequilibrium transport by a uniform DC electric field on a one-dimensional insulating tight-binding chain with the on-site disorder, using a finite-lattice calculation and the coherent potential approximation. We develop a theory of electric-field-assisted variable range hopping as a mechanism for nonlinear transport in a disordered chain. Our finite-lattice calculations of the electron propagation distance and the electron mobility determine the range of the variable range hopping as $\Delta<W\lesssim 2\Delta$ in the gap $\Delta$. We further propose a nonlinear scaling of the conductivity by an electric field that is similar to Mott’s variable range hopping in equilibrium. The nonlinear conductivity of an electronic lattice model follows the scaling law $\sigma(E)\propto\exp[-(E_{0}/E)^{\nu}]$ with the exponent $\nu=1/3$ in one dimension for the VRH. We also discuss the experimental relevance of temperature-dependent nonlinear current-voltage relation.
@@ I Introduction
The problem of disorder in solid-state systems has long been a significant area of focus in condensed matter physics. Anderson’s pioneering work [1] demonstrated that in a lattice with disordered potentials, electrons become localized in certain regions of the lattice, known as the Anderson localization (AL). This phenomenon has been studied extensively in the context of electronic systems in equilibrium [2, 3, 4, 5, 6, 7, 8], and the concept has since been expanded to encompass various wave phenomena [9, 10, 11, 12].
@@
with $d$ the dimensionality of the system by optimizing the Miller-Abraham’s formula for the hopping rate ${\cal W}$ at temperature $T$ between levels separated by distance $R$ and energy excitation by $\Delta\epsilon$,
@@
The purpose of this work is to deepen the theoretical understanding of the VRH in the nonlinear transport regime. While the role of strong fields in condensed matter is receiving a great deal of interest lately in nanoscale device applications, the theoretical framework has only been developed in recent decades, in particular, for quantum transport far from equilibrium. The problem is many-fold: theoretically, anything outside equilibrium cannot rely on textbook statistical mechanical principles such as Gibb’s ensemble, leaving the theory only with quite formidable options such as the Keldysh diagrammatic theory in such a regime [25, 26]. From the phenomenological perspective, nonequilibrium has quite a distinct reference state from an equilibrium that has entangled quantum and classical excitations that are often hard to distinguish. In the past, ad hoc scaling relations such as those obtained by replacing thermal energy with field-driven energy in thermodynamics relations [27] have often been used. In this work, we instead treat the disordered nonequilibrium limit rigorously by considering the electronic lattice model by directly solving the Keldysh theory. We test some notable results [28, 29, 30, 31, 32, 33, 34] that discuss the transport of electrons in disordered media in an arbitrary electric field.
@@
In our previous work [35], we addressed the spectral properties of a field-driven tight-binding model and validated Mott’s scaling of the VRH through Keldysh Green’s function method. The calculations highlighted that the electric field results in delocalization from the Anderson localization to a crossover to the Wannier Stark effect at very strong electric field, by using the spectral properties of the system. We also observed that in equilibrium the finite-lattice calculations showed exponential Lifshitz tails which are further broadened by the effect of the electric field and it was argued that this is crucial for understanding the transport behavior of disordered chains.
@@
In the present paper, we focus on the electronic transport of a disordered strongly-driven electron lattice. We present a simple Anderson model of disorder consisting of a tight-binding chain and on-site disordered potential and apply a DC electric field. We couple the chain to an infinite fermionic reservoir at each site, which ensures a steady-state solution through dissipation. We compute the nonequilibrium Green’s functions in the steady-state limit and compute the transport quantities. We explicitly demonstrate how the hopping mechanism evolves with an electric field in and out of the VRH limit by studying the propagation length. In nonequilibrium, we derive Mott’s scaling law as a function of the electric field, which we corroborate with Keldysh Green’s function theory. We also investigate the $IV$ (current-voltage) evolution in the VRH regime for nontrivial crossover from the hopping to Ohmic regime, which can be readily compared for nonlinear transport in quasi-one-dimensional systems. We finally discuss experimentally relevant temperature dependence of the field-activation of nonlinear conductance.
@@
Figure 1: Infinite tight-binding chain set-up with the electric field. The disorder-active region (red circles) of the chain consists of $N=501$ sites with Anderson disorder in the range $(-W,W)$. On both ends of the active region we have a semi-infinite chain without disorder that acts as leads (blue circles). Every site is coupled to a fermionic chain reservoir which helps maintain the steady-state limit out of equilibrium. Electrostatic potential $-lEa$ is applied across the entire chain including the leads.
@@
The rest of the paper is summarized as follows. In section-II, we summarize our quantum mechanical model introduced in our previous study [35] and discuss our calculations of Green’s functions, also discussed in great detail in the first paper, as well as transport properties. In section-III, we discuss our results for the disordered lattice case in equilibrium and nonequilibrium. We highlight some key aspects of our calculation against the approach of coherent potential approximation (CPA) for the disordered potential. Finally in section-IV, we discuss our results and conclude.
@@ II Disordered Chain under DC bias
Our model [35] consists of an infinite tight-binding chain, with a constant electric field applied to all sites and the central region (red circles) as the disorder-active region, as shown in Fig. 1. Each site at position $l$ in the main chain is at an electrostatic potential $-lEa$, $a$ being the lattice constant (set to unit length). With the setup, physical (gauge-covariant) observables such as electron density and current are translationally invariant in the zero disorder limit without any scattering between the disorder-active and the lead regions. Each site of the chain is coupled to fermion chains which act as a reservoir [36] as shown in Fig. 1. The Hamiltonian hence can be written as
@@
$\epsilon_{l}=2t+\Delta+V_{l}-lE$ (4)
where $E$ denotes the uniform DC electric field, $\Delta$ is the gap and $V_{l}$ is the on-site random disordered potential. The bath spectrum $\epsilon_{l\alpha}$ is similarly given as $\epsilon_{\alpha}-lE$ with the identical spectra $\epsilon_{\alpha}$ for every site $l$. Our system is considered to be a discretized limit of an insulator with the dispersion relation $\epsilon_{p}=p^{2}/2m+\Delta$ (with $11/(2m)=t=$ 1), that is displaced from the Fermi level as shown in Fig. 1. We set the gap $\Delta=0.3$ throughout our calculations.
@@
The reservoir consists of a fermion chain of length $L$ with identical dispersion $\epsilon_{\alpha}$ for each site. Its density of states is assumed structureless, constant, and with an infinite bandwidth. Mixing of the main orbitals with the reservoir is given by the hybridization function $\Gamma(\omega)=\pi g^{2}L^{-1}\sum_{\alpha}\delta(\omega-\epsilon_{\alpha})\equiv\Gamma$ in the infinite bandwidth limit. This parameter $\Gamma$, also referred to as dissipation rate in the text, denotes the energy level broadening and dephasing rate of electrons. The reservoir is kept at a temperature $T$ with Fermi level shifted by the electrostaic potential $-lE$. Hence the electron distribution at each site $l$ is given by the Fermi-Dirac function as $f_{0}(\omega-lE)=[1+e^{(\omega-lE)/T}]^{-1}$ in the reservoir. We incorporate the dissipation effects of the reservoir exactly by using the Keldysh Green’s function method and analyze its effect on the transport properties in the subsequent sections.
@@ II.1 Lattice Calculation of the conductance
We numerically calculate the full retarded Green’s function of the system for each disordered configuration. The general form of the $N\times N$ retarded Green’s function matrix is given by
@@
In Eq. (II.1), the term $-i\Gamma$ is the retarded reservoir self-energy causing broadening of levels and $t^{2}F^{R}_{-}$ and $t^{2}F^{R}_{+}$ are the self-energy terms of the semi-infinite leads attached on the left and right side of the disordered active sites, respectively, as depicted in Fig. 1. The retarded Green’s function of the right/left chains $F^{R}_{\pm}$, evaluated at the terminating orbital of the semi-infinite chain, are computed iteratively [37] as
$[F^{R}_{\pm}(\omega)]^{-1}=\omega+i\Gamma-(2t+\Delta)-t^{2}F^{R}_{\pm}(\omega\pm E).$ (8)
The Green’s functions on the disorder-active sites are obtained by matrix inversion using sparse matrix routines. In Eq. (7) the term $\Sigma^{<}_{k}(\omega)=2i\Gamma f_{0}(\omega+kE)+t^{2}\delta_{k,-N/2}F^{<}_{-}(\omega-NE/2)+t^{2}\delta_{k,N/2}F^{<}_{+}(\omega+NE/2)$ denotes the lesser self-energy at site $k$. Similarly, the lesser Green’s functions $F^{<}_{\pm}$ for the semi-infinite chains are computed iteratively as
$F^{<}_{\pm}(\omega)=|F^{R}_{\pm}(\omega)|^{2}[2i\Gamma f_{0}(\omega)+t^{2}F^{<}_{\pm}(\omega\pm E)]$ (9)
@@
[…]
@@
In the studies of the electron transport in disordered media [23], the Miller-Abraham’s relation, Eq. (2), has been the cornerstone of statistical theories, including Mott’s. Our approach puts the quantum mechanical problem by including the effects of electric field, disorder and dissipation in a first-principles manner and solve the Schrödinger equation of an electronic lattice directly by the Keldysh Green’s function method. More specifically, the first term in the argument of Eq. (2) for the wavefunction overlap is accounted for by the retarded Green’s functions $|{\cal G}^{R}_{ij}(\omega)|^{2}$ in Eq. (7), and the thermal factor in Eq. (2) is represented by the lesser quantities ${\cal G}^{<}_{ij}(\omega)$ or $\Sigma^{<}_{k}(\omega)$ in the framework of the Keldysh theory.
@@
As a benchmark, we complement the full-lattice calculation with the coherent potential approximation (CPA) [38, 39, 40] approach. This method is similar to the dynamical mean field theory [41, 25, 42] (DMFT) where the spatial inhomogeneity of the disordered lattice is replaced by a complex self-energy which represents the local effective medium potential. This is evaluated self consistently by embedding an impurity with a random potential drawn in the range of $-W\leq V\leq W$ and then computing the average Green’s function which is mapped to the effective medium Green’s function. We employ the nonequilibrium steady state calculation similar to the one discussed in the earlier works of one of the authors [37] along with the CPA approach to incorporate the effects of disorder to compare the lattice calculations. The CPA method is described in details in [35].
@@
Figure 2: (a) Spatial correlation function $C(x)$ plotted for different electric fields at disorder strength $W=0.4$, for $\Gamma=0.0005$. As the electric field $E$ increases, $C(x)$ widens, indicating the tendency of delocalization. As $E$ increases further, the signature of Bloch oscillations appears from large $x$. The range of Bloch oscillations scales well with $4t/E$ (as marked by the arrow). (b) Spatial correlation function at intermediate electric field ($E=0.03$) for different disorder values. The width of $C(x)$ reaches a maximum in the VRH regime. (c) Drift distance $L_{\rm drift}$ vs against electric field for disorder strengths [$W\in\{0.2,0.3,0.35,0.4,0.45,0.5,0.6,0.7,0.8\}$ in the order indicated by the arrow]. $L_{\rm drift}$ initially increases with $E$ due to the field-activation [43], and then decreases due to Bloch oscillations $x\sim 4t/E$ (marked by the dashed line). The initial rise at small fields is the strongest in the VRH regime. (d) Drift distance $L_{\rm drift}$ vs disorder strength at various electric fields ranging from $E=0.0025$ (dark blue) to $E=0.07$ (dark red) with a spacing of $\Delta E=0.0075$. The VRH range is demarcated by the black dashed lines.

### v3

@@ Abstract
Transport in disordered systems often occurs via the variable range hopping (VRH) in the dilute carrier density limit, where electrons hop between randomly distributed localized levels. We study the nonequilibrium transport by a uniform DC electric field on a one-dimensional insulating tight-binding chain with the on-site disorder, using a disordered-lattice calculation and the coherent potential approximation. We develop a theory of electric-field-assisted variable range hopping as a mechanism for nonlinear transport in a disordered chain. Our disordered-lattice calculations of the electron propagation distance and the electron mobility determine the range of the variable range hopping as $\Delta<W\lesssim 2\Delta$ in the gap $\Delta$. We further propose a nonlinear scaling of the conductivity by an electric field by extending Mott’s variable range hopping. The nonlinear conductivity of an electronic lattice model follows the scaling law $\sigma(E)\propto\exp[-(E_{0}/E)^{\nu}]$ with the exponent $\nu=1/3$ in one dimension for the VRH. We also discuss the experimental relevance of temperature-dependent nonlinear current-voltage relation.
@@ I Introduction
The problem of disorder in solid-state systems has long been a significant area of focus in condensed matter physics. Anderson’s pioneering work [1] demonstrated that in a lattice with disordered potentials, electrons become localized in certain regions of the lattice, known as the Anderson localization. This phenomenon has been studied extensively [2, 3, 4, 5, 6, 7, 8] in the context of electronic systems in equilibrium, and the concept has since been expanded to encompass various wave phenomena [9, 10, 11, 12].
@@
with $d$ the dimensionality of the system, by optimizing the Miller-Abraham’s formula for the hopping rate $\cal W$} at temperature $T$ between levels separated by distance $R$ and energy excitation by $\Delta\epsilon$,
@@
The purpose of this work is to deepen the theoretical understanding of the VRH in the nonlinear transport regime in an electronic lattice where the electron dynamics is explicitly considered. While the role of strong fields in condensed matter is receiving a great deal of interest lately in nanoscale device applications, the theoretical framework has only been developed in recent decades, particularly for quantum transport far from equilibrium. The problem is many-fold: theoretically, anything outside equilibrium cannot rely on textbook statistical mechanical principles such as Gibbs’ ensemble, leaving the theory only with quite formidable options such as the Keldysh diagrammatic theory in such a regime [24, 25]. From the phenomenological perspective, nonequilibrium has quite a distinct reference state from an equilibrium that has entangled quantum and classical excitations that are often hard to distinguish. In the past, ad hoc scaling relations such as those obtained by replacing thermal energy with field-driven energy in thermodynamics relations [26] have often been used. In this work, we instead treat the disordered nonequilibrium limit rigorously by considering the electronic lattice model by solving the Keldysh theory in a straightforward manner with the disorder averaging coming only at the end of the calculation. Therefore, we make a definitive statement for the scaling behavior of high-field conductivity in one of the most basic nonequilibrium models that contains the following minimum essential ingredients: lattice, Fermi-statistics, electric field, disorder, and dissipation. We test some notable results [27, 28, 29, 30, 31, 32, 33] that discuss the transport of electrons in disordered media in an arbitrary electric field.
@@
In our previous work [34], we addressed the spectral properties of a field-driven tight-binding model and validated Mott’s scaling of the VRH through Keldysh Green’s function method. The calculations highlighted that the electric field results in delocalization from the Anderson localization to a crossover to the Wannier Stark effect at a very strong electric field, by using the spectral properties of the system. We also observed that in equilibrium, the disordered-lattice calculations showed exponential Lifshitz tails, which are further broadened by the effect of the electric field, and it was argued that this is crucial for understanding the transport behavior of disordered chains. Furthermore, lattice-explicit calculations have revealed that the nonequilibrium distributions in the steady-state show highly nontrivial non-thermal evolution.
@@
In the present paper, we focus on the electronic transport of a disordered strongly-driven electron lattice. We present a simple Anderson model of disorder consisting of a tight-binding chain and an on-site disordered potential and apply a DC electric field. We couple the chain to an infinite fermionic reservoir at each site, which ensures a steady-state solution through dissipation. We compute the nonequilibrium Green’s functions in the steady-state limit and compute the transport quantities. We explicitly demonstrate how the hopping mechanism evolves with an electric field in and out of the VRH limit by studying the propagation length. In nonequilibrium, we derive Mott’s scaling law as a function of the electric field, which we corroborate with Keldysh Green’s function theory. We also investigate the $IV$ (current-voltage) evolution in the VRH regime for nontrivial crossover from the hopping to Ohmic regime, which can be readily compared for nonlinear transport in quasi-one-dimensional systems. We finally discuss experimentally relevant temperature dependence of the field-activation of nonlinear conductance.
@@
Figure 1: Infinite tight-binding chain set-up with the electric field. The disorder-active region (red circles) of the chain consists of $N=501$ sites with Anderson disorder in the range (W,W)(-W,W). On both ends of the active region, we have a semi-infinite chain without disorder that acts as leads (blue circles). Every site is coupled to a fermionic chain reservoir, which helps maintain the steady-state limit out of equilibrium. Electrostatic potential lEa-lEa is applied across the entire chain, including the leads and the reservoir levels. The main chain is kept above the Fermi level, marked by the shaded region by a gap $\Delta$.
@@
The rest of the paper is summarized as follows. In section-II, we summarize our quantum mechanical model introduced in our previous study [34] and discuss our calculations of Green’s functions, also discussed in great detail in the first paper, as well as transport properties. In section-III, we discuss our results for the disordered lattice case in equilibrium and nonequilibrium. We highlight some key aspects of our calculation against the approach of coherent potential approximation (CPA) for the disordered potential. Finally, in section-IV, we discuss our results and conclude.
@@ II Disordered Chain under DC bias
Our model [34] consists of an infinite tight-binding chain, with a constant electric field applied to all sites and the central region (red circles) as the disorder-active region, as shown in Fig. 1. Each site at position $l$ in the main chain is at an electrostatic potential lEa-lEa, $a$ being the lattice constant (set to unit length). With the setup, physical (gauge-covariant) observables such as electron density and current are translationally invariant in the zero disorder limit without any scattering between the disorder-active and the lead regions. Each site of the chain is coupled to fermion chains which act as a reservoir [35] as shown in Fig. 1. The Hamiltonian hence can be written as
@@
$\epsilon_{l}=2t+\Delta+V_{l}-l\cdot eEa$ (4)
where $E$ denotes the uniform DC electric field, $\Delta$ is the gap and $V_{l}$ is the on-site random disordered potential. The elecric charge $e$ and the lattice constant $a$ are set to 1, for the unit of charge and length, respectively. The bath spectrum $\epsilon_{l\alpha}$ is similarly given as $\epsilon_{\alpha}-lE$ with the identical spectra $\epsilon_{\alpha}$ for every site $l$. Our system is considered to be a discretized limit of an insulator with the dispersion relation $\epsilon_{p}=p^{2}/2m+\Delta$ (with $1/(2m)=t=1$), that is displaced from the Fermi level as shown in Fig. 1. We set the gap $\Delta=0.3$ throughout our calculations.
@@
The reservoir consists of a fermion chain of length $L$ with identical dispersion $\epsilon_{\alpha}$ for each site. Its density of states is assumed to be structureless, constant, and with an infinite bandwidth. Mixing of the main orbitals with the reservoir is given by the hybridization function $\Gamma(\omega)=\pi g^{2}L^{-1}\sum_{\alpha}\delta(\omega-\epsilon_{\alpha})\equiv\Gamma$ in the infinite bandwidth limit. This parameter $\Gamma$, also referred to as dissipation rate in the text, denotes the energy level broadening and dephasing rate of electrons. The reservoir is kept at a temperature $T$ with the Fermi level shifted by the electrostatic potential lE-lE. Hence the electron distribution at each site $l$ is given by the Fermi-Dirac function as $f_{\rm FD}(\omega-lE)=[1+e^{(\omega-lE)/T}]^{-1}$ in the reservoir. (We set the Boltzmann constant $k_{B}=1$.) We incorporate the dissipation effects of the reservoir exactly by using the Keldysh Green’s function method and analyze its effect on the transport properties in the subsequent sections.
@@ II.1 Lattice Calculation of the conductance
We numerically calculate the full retarded Green’s function of the system for each disordered configuration. The Green’s function of an infinite chain can be written in terms of the self-energy from the disorder-free regions. More detailed discussions on the method can be found in our previous work [34]. The general form of the $N\times N$ retarded Green’s function matrix is given by
@@
In Eq. (II.1), the term $i\Gamma$ is the retarded reservoir self-energy causing broadening of levels and $t^{2}F^{R}_{-}$ and $t^{2}F^{R}_{+}$ are the self-energy terms of the semi-infinite leads attached on the left and right side of the disordered active sites, respectively, as depicted in Fig. 1. The retarded Green’s function of the right/left chains $F^{R}_{\pm}$, evaluated at the terminating orbital of the semi-infinite chain, are computed iteratively [36] as
$F^{R}_{\pm}(\omega)]^{-1}=\omega+i\Gamma-(2t+\Delta)-t^{2}F^{R}_{\pm}(\omega\pm E).$ (8)
The Green’s functions on the disorder-active sites are obtained by matrix inversion using sparse matrix routines. In Eq. (7) the term $\Sigma^{<}_{k}(\omega)=2i\Gamma f_{\rm FD}(\omega+kE)+t^{2}\delta_{k,-N/2}F^{<}_{-}(\omega-NE/2)+t^{2}\delta_{k,N/2}F^{<}_{+}(\omega+NE/2)$ denotes the lesser self-energy at site $k$. Similarly, the lesser Green’s functions $F^{<}_{\pm}$ for the semi-infinite chains are computed iteratively as
$F^{<}_{\pm}(\omega)=|F^{R}_{\pm}(\omega)|^{2}[2i\Gamma f_{\rm FD}(\omega)+t^{2}F^{<}_{\pm}(\omega\pm E)]$ (9)
@@
The following parameters are tuned - disorder strength $W$, electric field $E$, temperature $T$, and dissipation $\Gamma$ to study transport quantities such as the local conductivity defined as :
$\sigma=\frac{\langle J_{l}\rangle}{E}$ (12)
and the mobility is defined as
$\mu=\frac{\sigma}{\langle n_{l}\rangle}$ (13)
where for simplicity $\langle f\rangle=\sum_{i}[\int dv_{i}P(v_{i})f_{i}]$ denote both averaging over spatial sites and disorder configurations. Similar to [34], we explore the electric field range in $E=10^{-4}$ to $0.1$ which corresponds to $10$ to $10^{4}\text{ kV/cm}$ that includes typically experimental scales. The dissipation parameters are considered smaller than lattice parameters and we keep $\Gamma$ in the order of $\Gamma\sim 10^{-4}$.
@@
In the studies of the electron transport in disordered media [6], the Miller-Abraham’s relation, Eq. (2), has been the cornerstone of statistical theories, including Mott’s. Our approach puts the quantum mechanical problem by including the effects of electric field, disorder, and dissipation in a first-principles manner and solves the Schrödinger equation of an electronic lattice directly by the Keldysh Green’s function method. More specifically, the first term in the argument of Eq. (2) for the wavefunction overlap is accounted for by the retarded Green’s functions $|{\cal G}^{R}_{ij}(\omega)|^{2}$ in Eq. (7), and the thermal factor in Eq. (2) is represented by the lesser quantities ${\cal G}^{<}_{ij}(\omega)$ or $\Sigma^{<}_{k}(\omega)$ in the framework of the Keldysh theory.
@@
As a benchmark, we complement the full-lattice calculation with the coherent potential approximation (CPA) [37, 38, 39] approach. This method is similar to the dynamical mean field theory [40, 24, 41] (DMFT) where the spatial inhomogeneity of the disordered lattice is replaced by a complex self-energy which represents the local effective medium potential. This is evaluated self-consistently by embedding an impurity with a random potential drawn in the range of $W\leq V\leq W$ and then computing the average Green’s function, which is mapped to the effective medium Green’s function. We employ the nonequilibrium steady-state calculation similar to the one discussed in the earlier works of one of the authors [36], along with the CPA approach to incorporate the effects of disorder to compare the disordered-lattice calculations. The CPA method is described in detail in [34]. The transport variables like the charge occupation and the conductivity are computed in the same manner as the disordered-lattice calculations, although they are only defined locally at the impurity site.
@@
Figure 2: (a) Spatial correlation function $C(x)$ plotted for different electric fields at disorder strength $W=0.4$, for $\Gamma=0.0005$. As the electric field $E$ increases, $C(x)$ widens, indicating the tendency of delocalization. As $E$ increases further, the signature of Bloch oscillations appears from large $x$. The range of Bloch oscillations scales well with $4t/E$ (as marked by the arrow). (b) Spatial correlation function at intermediate electric field ($E=0.03$) for different disorder values. The width of $C(x)$ reaches a maximum in the VRH regime. (c) Drift distance $L_{\rm drift}$ vs against electric field for disorder strengths $W\in\{0.2,0.3,0.35,0.4,0.45,0.5,0.6,0.7,0.8\}$ in the order indicated by the arrow]. $L_{\rm drift}$ initially increases with $E$ due to the field-activation [42], and then decreases due to Bloch oscillations $x\sim 4t/E$ (marked by the dashed line). The initial rise at small fields is the strongest in the VRH regime. (d) Drift distance $L_{\rm drift}$ vs disorder strength at various electric fields ranging from $E=0.0025$ (dark blue) to $E=0.07$ (dark red) with a spacing of $\Delta E=0.0075$. The VRH range is demarcated by the black dashed lines.